OpenStudy (cherie_magee):

Solve the following system of equations. 3x + 2y - 5 = 0 x = y + 10

OpenStudy (souvik):

put x=y+10 into 1st equation...

OpenStudy (anonymous):

given x=y+10 so in another equation if we replace the value of x we get 3(y+10)+2y-5=0 =>3y+30+2y-5=0 =>5y+25=0 =>5y=-25 =>y=-5 now in another equation we are gonna replace the value of y x=-5+10 =>x=5